Suppose the demand of jeans is given by P=21-5Qd and the supply is given by P=4+2Qs. Calculate the equilibrium quantity.


1
Expert's answer
2021-06-24T10:11:55-0400

"Qd=21-P\\Qs=P-4\\Qd=Qs\\21-P=P-4\\21+4=P+P\\25=2P\\P=\frac{25}{2}\\P=12.5\\Qd=21-P=21-12.5=8.5\\Qs=P-4=12.5-4=8.5"


The equilibrium quantity is therefore 8.5


Since quantity cannot be a decimal the equilibrium quantity is approximately 8